How much energy is consumed by a 1.00 Kilowatt hair dryer used for 13.0min and a 11.0 Watt night light that was left on for 21.0 hours?
Answer #1

a) E = P*t

= 1000*13*60

= 780000 J

b) E = P*t

= 11*21*60*60

= 831600 J
Answer #2

apply Power P = Eenrgy./timr

so E = Pt

E of hair drier = 1000 * 13*60 = 780 K Joules
----------------------

E of light = 11* 21* 60 * 60 = 831.6 k Joules
Answer #3

Energy by hair dryer = 1000* 13/60 =216.67 Wh

Energy by bulb = 11*21 =231 Wh

Total energy = 216.67+231=447.67 Wh = 0.45 kWh (approx)